OpenStudy (kay.baby.):

WILL MEDAL Which rational number equals 0 point 3 with bar over 3? 1 over 5 1 over 3 3 over 10 3 over 5

OpenStudy (kay.baby.):

My guess would be 3 over 10

OpenStudy (danieldjpon3):

ya

OpenStudy (danieldjpon3):

.3 = 3 tenths

OpenStudy (danieldjpon3):

.3 = 3 tenths

OpenStudy (danieldjpon3):

oops

OpenStudy (kay.baby.):

thanks can you help with one more

OpenStudy (danieldjpon3):

kk

OpenStudy (danieldjpon3):

sure

OpenStudy (kay.baby.):

If (82)p = 84, what is the value of p? 2 3 4 6

OpenStudy (danieldjpon3):

ok

OpenStudy (kay.baby.):

(8^2)^p = 8^4

OpenStudy (danieldjpon3):

i think ikonw this one

OpenStudy (kay.baby.):

Okay :)

OpenStudy (danieldjpon3):

isnt it (8^2)p = 84?

OpenStudy (danieldjpon3):

check if it is

OpenStudy (danieldjpon3):

bc i've done one like the one I said

OpenStudy (kay.baby.):

(8^2) p = 8^4

OpenStudy (danieldjpon3):

ok

OpenStudy (danieldjpon3):

ez ive done one like this one

OpenStudy (danieldjpon3):

16 times p

OpenStudy (danieldjpon3):

then 8x8x8x8 is 1024

OpenStudy (danieldjpon3):

so 1024 divided by 16

OpenStudy (danieldjpon3):

equals p

OpenStudy (danieldjpon3):

answer is 64

OpenStudy (kay.baby.):

my options are 2, 3 4 or 6

OpenStudy (danieldjpon3):

hmm

OpenStudy (danieldjpon3):

send me screen shot of ur question

OpenStudy (kay.baby.):

i would say 2 because it would make them congruent

OpenStudy (kay.baby.):

okay one sec

OpenStudy (danieldjpon3):

(8^2) p = 8^4

OpenStudy (kay.baby.):

yes

OpenStudy (danieldjpon3):

but them if (8^2) times p = to 32

OpenStudy (danieldjpon3):

and 8^4 doesnt equal to 32

OpenStudy (danieldjpon3):

are you sure there's no negetives in this?

OpenStudy (kay.baby.):

no negatives

OpenStudy (danieldjpon3):

k

OpenStudy (danieldjpon3):

let's try 6

OpenStudy (danieldjpon3):

16 times 6 = 72

OpenStudy (danieldjpon3):

8 time 8 = 16 times 8 is 128 times 8 is 1024

OpenStudy (danieldjpon3):

Can you please send me the screen shot, i don't get it

OpenStudy (kay.baby.):

it does not allow me to screenshot im on my laptop

OpenStudy (danieldjpon3):

do you have a PrtSc button on your keyboard?

OpenStudy (danieldjpon3):

Is your laptop running on windows 10?

OpenStudy (kay.baby.):

yes

OpenStudy (danieldjpon3):

kk then you should have a print screen button

OpenStudy (danieldjpon3):

press your home buttton while pressing PrtSc

OpenStudy (kay.baby.):

okay:)

OpenStudy (danieldjpon3):

kk hopefully it works

OpenStudy (kay.baby.):

i got the screenshot but it wont upload thanks for all the help tho

OpenStudy (danieldjpon3):

kk

OpenStudy (danieldjpon3):

I still don't get it

OpenStudy (kay.baby.):

thats okay

OpenStudy (danieldjpon3):

weird

OpenStudy (danieldjpon3):

try to copy teh question and paste it in google and write open study after it

OpenStudy (kay.baby.):

kk thanks again